A ball is dropped from a building of height 45m.simultaneously another ball is thrown up with a speed 40m/s.the relative speeds of balls varies with time as

Grade:11

2 Answers

Vikas TU
6 years ago
Dear Student,
A ball is dropped from a working of stature 45m, at that point its speed is given by 
 
Let a be the increasing speed because of gravity 
 
Vb = at - (1.) 
 
Presently the at the same time another ball is hurled with speed of 40 m/sec 
 
v2b = 40 – at 
 
As the speed is inverse to each other, at that point these speed will be included for the relative speed. 
 
vb + v2b = at + 40 – at = 40 m/sec .
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)
Paravasthu Nava Deep
5 years ago
Va(velocity of first ball)= gt
Vb(velocity of the second ball)= 40m/s-gt
Rlative velocity of the ball wuth as = Vab = Va + Vb = gt+(40-gt)
=gt+40-gt= 40
gt-gt = 0
g(t-t) = 0
t-t = 0
t = t =>t/t = 1
There fore relative speed of the ball varies with time as t0 = 1